if 2 students move from van a to b the two van will have the same number of students but if two students move from van b to a van b will have half the number of students that were in a. Show at least one equation to get a mark.
Answer by JulietG(1812)

A - 2 (van A is losing 2 students) = B
B - 2 (van B losing 2 students) = .5A (half the amount of A)
A-2=B
B-2=.5A
Let's take the known value of B from the first equation and substitute it in the second.
B-2=.5A
(A-2)-2 = .5A
A-4 = .5A
Add 4 to each side
A = .5A + 4
If half of A + 4 is equal to A, then A must be twice 4.
A = 8
If A equals 8, then B = 6